# Levi-Civita connection on the irreducible quantum flag manifolds

### v1

@@ 5.3. The opposite complex structure on the calculus
Suppose $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ is a factorizable $A$-covariant complex structure on a quantum homogeneous space $B=A^{\mathrm{co}(H)}$, where $H$ is a compact quantum group algebra and $\Omega^{1}$ is an object of ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Let $\mathscr{H}_{1}$ and $\mathscr{H}_{2}$ be the $A$-covariant Hermitian metrics on $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ constructed in Theorem 4.4. Then the connection $\nabla=\nabla_{\mathrm{Ch}}+\nabla_{\mathrm{Ch},\mathrm{op}}$ of Theorem 5.8 is an $A$-covariant bimodule connection on $\Omega^{1}$.
@@ 6.1. Drinfeld–Jimbo quantum groups
Let $\mathfrak{g}$ be a finite-dimensional complex simple Lie algebra of rank $r$. We fix a Cartan subalgebra $\mathfrak{h}$ with corresponding root system $\Delta\subseteq\mathfrak{h}^{\ast}$, where $\mathfrak{h}^{\ast}$ denotes the linear dual of $\mathfrak{h}$. With respect to a choice of simple roots $\Pi=\{\alpha_{1},\alpha_{2},\cdots\alpha_{r}\}$, denote by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) the symmetric bilinear form induced on $\mathfrak{h}^{\ast}$ by the Killing form of $\mathfrak{g}$, normalised so that any shortest simple root $\alpha_{i}$ satisfies $(\alpha_{i},\alpha_{i})=2$. The coroot $\alpha_{i}^{\vee}$ of a simple root $\alpha_{i}$ is defined by
@@
Let $q\in\mathbb{R}$ such that $q\notin\{-1,0,1\}$, and denote $q_{i}:=q^{d_{i}}$. The quantised universal enveloping algebra $U_{q}(\mathfrak{g})$ is the noncommutative associative algebra generated by the elements $E_{i},F_{i},K_{i}$, and $K_{i}^{-1}$, for $i=1,\cdots,r$, satisfying the relations (12) - (16) of [KS, Chapter 6].
@@ A Hopf $\ast$-algebra structure, called the compact real form of $U_{q}(\mathfrak{g})$, is defined by
Moreover, $v_{\mu}$ is the unique such element up to scalar multiple. We call any finite direct sum of such $U_{q}(\mathfrak{g})$-representations a finite-dimensional type-$1$ representation. For further details on Drinfeld–Jimbo quantised enveloping algebras, we refer the reader to the standard texts [CP, KS], or to the seminal papers [Dri, Jim].
@@ 6.2. Quantum Coordinate Algebras and irreducible flag manifolds
In this subsection we recall some necessary material about quantised coordinate algebras. Let $V$ be a finite-dimensional left $U_{q}(\mathfrak{g})$-module, $v\in V$, and $f\in V^{*}$, the $\mathbb{C}$-linear dual of $V$, endowed with its right $U_{q}(\mathfrak{g})$-module structure.
@@ 6.3. Some consequences of Takeuchi’s equivalence
We recall that an element $X$ of the centre $Z(U_{q}(\mathfrak{l}_{S}))$ of $U_{q}(\mathfrak{l}_{S})$ acts on any irreducible finite-dimensional type-1 $U_{q}(\mathfrak{l}_{S})$-module $W$ by a central character $\chi_{W}\in\operatorname{Hom}(Z(U_{q}(\mathfrak{l}_{S})),\mathbb{C})$, i.e,
@@
Now, it was noted in [DKÓSS2, §4.4] (also see the discussion in [DKÓSS1, §4.5]) that there is a distinguished group-like element $Z$ in the centre of $U_{q}(\mathfrak{l}_{S})$ defined as
@@
In fact, from the proof of Theorem 4.12 of [DKÓSS2], we know that
@@
Since $(\varpi_{s},\alpha_{s})\neq 0$ and the elements of the centre of $U_{q}(\mathfrak{l}_{S})$ separate irreducible finite-dimensional type-1 $U_{q}(\mathfrak{l}_{S})$-modules, we have that (37a) holds. The proofs of the equivalences (37b) and (37c) are similar. ∎
@@
It is easy to show that $\dim\left(V^{(1,1)}\right)^{U_{q}(\mathfrak{l}_{S})}=1$, for example, see [DKÓSS2, Lemma 3.4]. Therefore, there is $\lambda\in\mathbb{C}^{\times}$ such that $\Phi(\wedge\circ\operatorname{coev}_{1})(1)=\lambda\Phi(\wedge\circ\operatorname{coev}_{2})(1)$ which proves the claim. ∎
@@ 6.4. Metrics on the Heckenberger–Kolb calculi
We start by observing that $(g^{1,0})^{\dagger}=g^{1,0}$ and $(g^{0,1})^{\dagger}=g^{0,1}$. Indeed, this follows by applying a similar proof to that of Corollary 2.13 and using the equations (26) and (27). Thus,
@@
Thus, $g=g^{1,0}-\lambda g^{0,1}$, where $\lambda\in\mathbb{R}\setminus\{0\}.$ Hence, by Theorem 3.6, we conclude that $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a real $\mathcal{O}_{q}(G)$-covariant metric. Since $\wedge g^{1,0}=\lambda\wedge g^{0,1}$, $g$ is quantum symmetric. ∎
@@
The set of all covariant metrics on $\Omega^{1}_{q}(G/L_{S})$ is a two parameter family. Moreover, there exists a unique (up to scalar) quantum symmetric $\mathcal{O}_{q}(G)$-covariant metric on $\Omega^{1}_{q}(G/L_{S})$.
@@
We have already proved the existence of a covariant real quantum symmetric metric $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ in Theorem 6.10, where $g=g^{1,0}-\lambda g^{0,1}$, where $\lambda\in\mathbb{R}\setminus\{0\}$. Let $(\widetilde{g},\widetilde{(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )})$ be another covariant quantum symmetric metric. Then by Theorem 3.2, $\widetilde{g}=\sigma.g$ for an automorphism $\sigma=(\lambda_{1},\lambda_{2})\in\mathbb{C}^{\times}\times\mathbb{C}^{\times}$ of $V=V^{(1,0)}\oplus V^{(0,1)}.$ Thus, by (6),
@@ 6.5. The Levi-Civita connection
By Remark 2.10, $g=\operatorname{coev}(1)$ for a left $A$-covariant map
$\operatorname{coev}:B\rightarrow\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S}).$
Therefore, $g$ is an $A$-coinvariant element of $\Omega^{1}\otimes_{B}\Omega^{1}$. Hence, $\nabla g$ is an $\mathcal{O}_{q}(G)$ coinvariant element of $\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S}).$ Therefore, $\nabla g=0$ by (36a) and so $\nabla$ is compatible with the metric $g$ in the sense of Definition 2.21.
@@
The uniqueness follows from the fact that there is a unique covariant connection on $\Omega^{1}_{q}(G/L_{S})$ and follows along the lines of Theorem 4.5 of [DKÓSS1]. Indeed, let $\nabla_{1}$ and $\nabla_{2}$ are two covariant connections on $\Omega^{1}_{q}(G/L_{S}).$ Then $\nabla_{1}-\nabla_{2}$ is a covariant and left $\mathcal{O}_{q}(G)$ linear map from $\Omega^{1}_{q}(G/L_{S})$ to $\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S})$ so that $\Phi(\nabla_{1}-\nabla_{2})\in\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V,V\otimes V).$
Since $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ are simple, the proof of Theorem 4.5 of [DKÓSS1] shows that $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(1,0)},V\otimes V^{(1,0)})$ and $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(0,1)},V\otimes V^{(0,1)})$ are zero. By a similar computation, $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(1,0)},V\otimes V^{(0,1)})$ and $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(0,1)},V\otimes V^{(1,0)})$ are also zero. Therefore, $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V,V\otimes V)=0$ implying that $\nabla_{1}-\nabla_{2}=0$. ∎
@@ 7.1. Some preparatory results
Let $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ be a complex structure on a $\ast$-algebra $B$ and $\mathcal{E}$ be a $B$-bimodule which is finitely generated as a left $B$-module and $\mathscr{H}$ be an Hermitian metric on $\mathcal{E}$. If $P$ denotes the matrix defined in (41), then with the notations as above, we have
@@ 7.2. The $(1,0)$-part of the Chern connection
Our construction of the $(1,0)$-part of the Chern connection is inspired by and closely related to the formula derived in [OBvvR, Proposition 7.3]. However, our novelty lies in the usage of the language of bar-categories which makes the construction more transparent.
@@
Suppose $\overline{\partial}_{\mathcal{E}}$ is a left $\overline{\partial}$-connection on a $B$-bimodule $\mathcal{E}$ as in Proposition 7.5 and $\widetilde{\overline{\partial}}_{\mathcal{E}}$ is the right $\overline{\partial}$-connection on $\prescript{}{B}{\operatorname{Hom}(\mathcal{E},B)}$ defined in (46) and $\mathscr{H}$ be an Hermitian metric on $\mathcal{E}$ as in Definition 4.1. Then the map ¯~,:¯¯${\widetilde{\overline{\partial}}}_{\mathcal{E},\mathscr{H}}:\overline{\mathcal{E}}\to\overline{\mathcal{E}}\otimes_{B}\Omega^{(0,1)}$ defined as
@@ 8.2. Regarding Takeuchi’s equivalence
$U^{-1}(\sum_{i}a_{i}\otimes[m_{i}])=\sum_{i}a_{i}S((m_{i})_{(-1)})(m_{i})_{(0)}.$ (60)
The aim of this subsection is to state Proposition 8.4. Through this result is well-known to the experts, we include its proof since it has been used in the article more than once. We will need to introduce some more notations.
@@
where we have used (59), (60), the fact that $U$ is a right $B$-module map and that the right $B$-module structure of $\Phi(M)$ is trivial.
@@
Suppose $B=A^{\mathrm{co}(H)}$ is a quantum homogeneous space equipped with an $A$-covariant differential calculus such that $\Omega^{1}$ is an object of the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Then $(\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1})^{\mathrm{co}(A)}=0$ if and only if the only morphism from $\mathbb{C}$ to $\Phi(\Omega^{1})\otimes\Phi(\Omega^{1})\otimes\Phi(\Omega^{1})$ in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ is the zero morphism.
@@
This follows by observing that there is a one to one correspondence between $(\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1})^{\mathrm{co}(A)}$ and the set of all left $B$-linear left $A$-colinear maps from $B$ to $\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1}$, defined by $X\mapsto f_{X}$, where $f_{X}:B\rightarrow\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1}$ is defined as $f_{X}(b)=bX.$ As $f_{X}$ is left $B$-linear and left $A$-colinear, Proposition 8.4 implies that $f_{X}$ is a morphism in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Now the result follows from Takeuchi’s equivalence. ∎

### v2

@@ 5.3. The opposite complex structure on the calculus
Suppose $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ is a factorizable $A$-covariant complex structure on a quantum homogeneous space $B=A^{\mathrm{co}(H)}$ satisfying the hypotheses of Theorem 4.4. Let $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ be a real $A$-covariant metric on $\Omega^{1}$ and $\mathscr{H}_{1},\mathscr{H}_{2}$ be the $A$-covariant Hermitian metrics on $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ constructed in Theorem 4.4. Then the connection $\nabla=\nabla_{\mathrm{Ch}}+\nabla_{\mathrm{Ch},\mathrm{op}}$ of Theorem 5.8 is an $A$-covariant bimodule connection on $\Omega^{1}$.
@@ 6.1. Drinfeld–Jimbo quantum groups
Let $\mathfrak{g}$ be a finite-dimensional complex simple Lie algebra of rank $r$. We fix a Cartan subalgebra $\mathfrak{h}$ with corresponding root system $\Delta\subseteq\mathfrak{h}^{\ast}$, where $\mathfrak{h}^{\ast}$ denotes the linear dual of $\mathfrak{h}$. With respect to a choice of simple roots $\Pi=\{\alpha_{1},\alpha_{2},\cdots\alpha_{r}\}$, denote by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) the symmetric bilinear form induced on $\mathfrak{h}^{\ast}$ by the Killing form of $\mathfrak{g}$, normalized so that any shortest simple root $\alpha_{i}$ satisfies $(\alpha_{i},\alpha_{i})=2$. The coroot $\alpha_{i}^{\vee}$ of a simple root $\alpha_{i}$ is defined by
@@
Let $q\in\mathbb{R}$ such that $q\notin\{-1,0,1\}$, and denote $q_{i}:=q^{d_{i}}$. The quantized universal enveloping algebra $U_{q}(\mathfrak{g})$ is the noncommutative associative algebra generated by the elements $E_{i},F_{i},K_{i}$, and $K_{i}^{-1}$, for $i=1,\cdots,r$, satisfying the relations (12) - (16) of [KS, Chapter 6].
@@ A Hopf $\ast$-algebra structure, called the compact real form of $U_{q}(\mathfrak{g})$, is defined by
Moreover, $v_{\mu}$ is the unique such element up to scalar multiple. We call any finite direct sum of such $U_{q}(\mathfrak{g})$-representations a finite-dimensional type-$1$ representation. For further details on Drinfeld–Jimbo quantized enveloping algebras, we refer the reader to the standard texts [CP, KS], or to the seminal papers [Dri, Jim].
@@ 6.2. Quantum Coordinate Algebras and irreducible flag manifolds
In this subsection we recall some necessary material about quantized coordinate algebras. Let $V$ be a finite-dimensional left $U_{q}(\mathfrak{g})$-module, $v\in V$, and $f\in V^{*}$, the $\mathbb{C}$-linear dual of $V$, endowed with its right $U_{q}(\mathfrak{g})$-module structure.
@@ 6.3. Some consequences of Takeuchi’s equivalence
We recall that an element $X$ of the center $Z(U_{q}(\mathfrak{l}_{S}))$ of $U_{q}(\mathfrak{l}_{S})$ acts on any irreducible finite-dimensional type-1 $U_{q}(\mathfrak{l}_{S})$-module $W$ by a central character $\chi_{W}\in\operatorname{Hom}(Z(U_{q}(\mathfrak{l}_{S})),\mathbb{C})$, i.e,
@@
Now, it was noted in [DKOSS2, §4.4] (also see the discussion in [DKOSS1, §4.5]) that there is a distinguished group-like element $Z$ in the center of $U_{q}(\mathfrak{l}_{S})$ defined as
@@
In fact, from the proof of Theorem 4.12 of [DKOSS2], we know that
@@
Since $(\varpi_{s},\alpha_{s})\neq 0$ and the elements of the center of $U_{q}(\mathfrak{l}_{S})$ separate irreducible finite-dimensional type-1 $U_{q}(\mathfrak{l}_{S})$-modules, we have that (37a) holds. The proofs of the equivalences (37b) and (37c) are similar. ∎
@@
It is easy to show that $\dim\left(V^{(1,1)}\right)^{U_{q}(\mathfrak{l}_{S})}=1$, for example, see [DKOSS2, Lemma 3.4]. Therefore, there is $\lambda\in\mathbb{C}^{\times}$ such that $\Phi(\wedge\circ\operatorname{coev}_{1})(1)=\lambda\Phi(\wedge\circ\operatorname{coev}_{2})(1)$ which proves the claim. ∎
@@ 6.4. Metrics on the Heckenberger–Kolb calculi
We start by observing that $(g^{1,0})^{\dagger}=g^{1,0}$ and $(g^{0,1})^{\dagger}=g^{0,1}$. Indeed, let us define
$\xi^{1,0}:\Omega^{(1,0)}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{(0,1)}_{q}(G/L_{S})\to\mathcal{O}_{q}(G/L_{S})$
as in the proof of Theorem 3.6. Then by (26), we have
$(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0}=\xi^{1,0}.$ (39)
[…]
It can be easily checked that $(\Omega^{(0,1)}_{q}(G/L_{S}),\xi^{1,0},\eta)$ is a right dual of $\Omega^{(1,0)}_{q}(G/L_{S}).$ Since $(\Omega^{(0,1)}_{q}(G/L_{S}),(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )^{1,0},\operatorname{coev}^{1,0})$ is also a right dual of $\Omega^{(1,0)}_{q}(G/L_{S}),$ (39) implies that $\operatorname{coev}^{1,0}=\eta,$ i.e, $g^{1,0}=(g^{1,0})^{\dagger}.$
Similarly, it follows that $(g^{0,1})^{\dagger}=g^{0,1}$.
Thus,
@@
Thus, $g=g^{1,0}-\lambda g^{0,1}$, where $\lambda\in\mathbb{R}\setminus\{0\}.$ Hence, by Theorem 3.6, we conclude that $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is a real $\mathcal{O}_{q}(G)$-covariant metric. Since $\wedge g^{1,0}=\lambda\wedge g^{0,1}$, $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ is quantum symmetric. ∎
@@
The set of all $\mathcal{O}_{q}(G)$-covariant metrics on $\Omega^{1}_{q}(G/L_{S})$ is a two parameter family. Moreover, there exists a unique (up to scalar) quantum symmetric $\mathcal{O}_{q}(G)$-covariant metric on $\Omega^{1}_{q}(G/L_{S})$.
@@
We have already proved the existence of a covariant real quantum symmetric metric $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ in Theorem 6.10, where $g=g^{1,0}-\lambda g^{0,1}$, where $\lambda\in\mathbb{R}\setminus\{0\}$. Let $(\widetilde{g},\widetilde{(\leavevmode\nobreak\ ,\leavevmode\nobreak\ )})$ be another covariant quantum symmetric metric. Then by Theorem 3.2, $\widetilde{g}=\tau.g$ for an automorphism $\tau=(\lambda_{1},\lambda_{2})\in\mathbb{C}^{\times}\times\mathbb{C}^{\times}$ of $V=V^{(1,0)}\oplus V^{(0,1)}.$ Thus, by (6),
@@ 6.5. The Levi-Civita connection
By Remark 2.10, $g=\operatorname{coev}(1)$ for a left $\mathcal{O}_{q}(G)$-covariant map
$\operatorname{coev}:\mathcal{O}_{q}(G/L_{S})\rightarrow\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S}).$
Therefore, $g$ is an $\mathcal{O}_{q}(G)$-coinvariant element of $\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S})$. Hence, $\nabla g$ is an $\mathcal{O}_{q}(G)$ coinvariant element of $\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S}).$ Therefore, $\nabla g=0$ by (36a) and so $\nabla$ is compatible with the metric $(g,(\leavevmode\nobreak\ ,\leavevmode\nobreak\ ))$ in the sense of Definition 2.21.
@@
The uniqueness follows from the fact that there is a unique covariant connection on $\Omega^{1}_{q}(G/L_{S})$ and follows along the lines of Theorem 4.5 of [DKOSS1]. Indeed, let $\nabla_{1}$ and $\nabla_{2}$ are two covariant connections on $\Omega^{1}_{q}(G/L_{S}).$ Then $\nabla_{1}-\nabla_{2}$ is a covariant and left $\mathcal{O}_{q}(G)$ linear map from $\Omega^{1}_{q}(G/L_{S})$ to $\Omega^{1}_{q}(G/L_{S})\otimes_{\mathcal{O}_{q}(G/L_{S})}\Omega^{1}_{q}(G/L_{S})$ so that $\Phi(\nabla_{1}-\nabla_{2})\in\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V,V\otimes V).$
Since $\Omega^{(1,0)}$ and $\Omega^{(0,1)}$ are simple, the proof of Theorem 4.5 of [DKOSS1] shows that $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(1,0)},V\otimes V^{(1,0)})$ and $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(0,1)},V\otimes V^{(0,1)})$ are zero. By a similar computation, $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(1,0)},V\otimes V^{(0,1)})$ and $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V^{(0,1)},V\otimes V^{(1,0)})$ are also zero. Therefore, $\operatorname{Hom}_{U_{q}(\mathfrak{l}_{S})}(V,V\otimes V)=0$ implying that $\nabla_{1}-\nabla_{2}=0$. ∎
@@ 7.1. Some preparatory results
Let $(\Omega^{(\bullet,\bullet)},\wedge,\partial,\overline{\partial})$ be a complex structure on a $\ast$-algebra $B$ and $\mathcal{E}$ be a $B$-bimodule which is finitely generated and projective as a left $B$-module and $\mathscr{H}$ be an Hermitian metric on $\mathcal{E}$. If $P$ denotes the matrix defined in (42), then with the notations as above, we have
@@ 7.2. The $(1,0)$-part of the Chern connection
Our construction of the $(1,0)$-part of the Chern connection is inspired by and closely related to the formula derived in [OSvR, Proposition 7.3]. However, our novelty lies in the usage of the language of bar-categories which makes the construction more transparent.
@@
Suppose $\overline{\partial}_{\mathcal{E}}$ is a left $\overline{\partial}$-connection on a $B$-bimodule $\mathcal{E}$ as in Proposition 7.5, $\widetilde{\overline{\partial}}_{\mathcal{E}}$ is the right $\overline{\partial}$-connection on $\prescript{}{B}{\operatorname{Hom}(\mathcal{E},B)}$ defined in (47) and $\mathscr{H}$ is an Hermitian metric on $\mathcal{E}$ as in Definition 4.1. Then the map ¯~,:¯¯${\widetilde{\overline{\partial}}}_{\mathcal{E},\mathscr{H}}:\overline{\mathcal{E}}\to\overline{\mathcal{E}}\otimes_{B}\Omega^{(0,1)}$ defined as
@@ 8.2. Regarding Takeuchi’s equivalence
$U^{-1}\left(\sum_{i}a_{i}\otimes[m_{i}]\right)=\sum_{i}a_{i}S((m_{i})_{(-1)})(m_{i})_{(0)}.$ (61)
The aim of this subsection is to state Proposition 8.4. Though this result is well-known to the experts, we include its proof since it has been used in the article more than once. We will need to introduce some more notations.
@@
where we have used (60), (61) and that the right $B$-module structure of $\Phi(M)$ is trivial.
@@
Suppose $B=A^{\mathrm{co}(H)}$ is a quantum homogeneous space equipped with an $A$-covariant differential calculus such that $\Omega^{1}$ is an object of the category ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Then (${}^{\mathrm{co}(A)}(\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1})=0$ if and only if the only morphism from $\mathbb{C}$ to $\Phi(\Omega^{1})\otimes\Phi(\Omega^{1})\otimes\Phi(\Omega^{1})$ in ${{}^{\phantom{H}\mathllap{H}}_{\phantom{H}\mathllap{}}\mathsf{mod}}$ is the zero morphism.
@@
This follows by observing that there is a one to one correspondence between (${}^{\mathrm{co}(A)}(\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1})$ and the set of all left $B$-linear left $A$-colinear maps from $B$ to $\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1}$, defined by $X\mapsto f_{X}$, where $f_{X}:B\rightarrow\Omega^{1}\otimes_{B}\Omega^{1}\otimes_{B}\Omega^{1}$ is defined as $f_{X}(b)=bX.$ As $f_{X}$ is left $B$-linear and left $A$-colinear, Proposition 8.4 implies that $f_{X}$ is a morphism in ${}^{\phantom{B}\mathllap{A}}_{\phantom{B}\mathllap{B}}\mathsf{mod}_{0}$. Now the result follows from Takeuchi’s equivalence. ∎